### v1

@@ Abstract
Hyperspectral images (HSIs) are often contaminated by a mixture of noises such as Gaussian noise, dead lines, stripes, and so on. In this paper, we propose a novel approach for HSI denoising and destriping, called NLTL2p, which consists of an orthogonal constrained minimization model and an iterative algorithm with convergence guarantees. The model of the proposed NLTL2p approach is built based on a new sparsity-enhanced Nonlocal Low-rank Tensor regularization and a tensor $\ell_{2,p}$ norm with $p\in(0,1)$. The low-rank constraints for HSI denoising utilize the spatial nonlocal self-similarity and spectral correlation of HSIs and are formulated based on independent higher-order singular value decomposition with sparsity enhancement on its core tensor to prompt more low-rankness. The tensor $\ell_{2,p}$ norm for HSI destriping is extended from the matrix $\ell_{2,p}$ norm. A proximal block coordinate descent algorithm is proposed in the NLTL2p approach to solve the resulting nonconvex nonsmooth minimization with orthogonal constraints. We show any accumulation point of the sequence generated by the proposed algorithm converges to a first-order stationary point, which is defined using three equalities of substationarity, symmetry, and feasibility for orthogonal constraints. In the numerical experiments, we compare the proposed method with state-of-the-art methods including a deep learning based method, and test the methods on both simulated and real HSI datasets. Our proposed NLTL2p method demonstrates outperformance in terms of metrics such as mean peak signal-to-noise ratio as well as visual quality.
@@ 1 Introduction
Hyperspectral images (HSIs) are collected by hyperspectral sensors across the electromagnetic spectrum. For a three-dimensional (3-D) HSI, the first two dimensions represent spatial information, and the third dimension represents the spectral information of a scene. An illustration of an HSI is shown in Fig. 1. HSIs are widely used for various applications [27, 29, 28, 46] such as object detection [39], material identification [6, 15], etc.
@@
Recently, some deep neural networks [4, 41, 34] have been proposed to denoise HSIs. Quasi-Recurrent Neural Networks (QRNN) [4] combined recurrent neural networks (RNN) with convolutional neural networks (CNN); the 3-D version of QRNN (QRNN3D) [34] can effectively embed the correlation of HSIs; CNN can also be used as a denoiser in a plug-and-play fashion for HSI denoising [30].
To remove Gaussian noise and stripes simultaneously, we propose an optimization model utilizing low-rank tensor regularization and a group sparsity measure, which is formulated as follows
(1) $\begin{split}\min_{\begin{subarray}{c}\mathcal{S},[{X}_{1}],[{X}_{2}],\\ [{X}_{3}],[\mathcal{G}],\mathcal{L}\end{subarray}}\quad&\frac{\delta}{2}\|\mathbf{R}(\mathcal{L}+\mathcal{S}-\mathcal{D})\|_{F}^{2}+\gamma\|\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{S}\|^{p}_{2,p}+\|[\mathcal{G}]\|_{1,w}\\ &+\frac{1}{2}\|\mathbf{R}(\mathcal{L})-[\mathcal{G}]\times_{1}[X_{1}]\times_{2}[X_{2}]\times_{3}[X_{3}]\|^{2}_{F}\\ \mbox{s.t. }&[{X}_{i}]^{\top}[{X}_{i}]=[\mathbf{I}_{n_{i}}],\quad i=1,2,3,\end{split}$
@@
  • $\mathbf{R}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$ denotes the similar blocks extraction operator, and $\mathbf{R}^{\top}:$ $\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}\to\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ denotes the transpose of $\mathbf{R}$ satisfying $\langle\mathbf{R}(\mathcal{L}),[\mathcal{Y}]\rangle=\langle\mathcal{L},\mathbf{R}^{\top}([\mathcal{Y}])\rangle$ for any $\mathcal{L}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, $[\mathcal{Y}]\in\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$;
  • $\mathcal{W}_{\mathbf{R}}\in\mathbb{R}_{++}^{I_{1}\times I_{2}\times I_{3}}$ denotes a weight tenor such that the component-wise multiplication with $\mathcal{W}_{\mathbf{R}}$ is an equivalent operation of $\mathbf{R}^{\top}\mathbf{R}$, i.e., $\mathcal{W}_{\mathbf{R}}\odot\operatorname{Id}=\mathbf{R}^{\top}\mathbf{R}$, $\odot$ denotes the component-wise multiplication, $\operatorname{Id}$ denotes the identity mapping on $\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and $\sqrt{\mathcal{W}_{\mathbf{R}}}$ denotes the component-wise square root of $\mathcal{W}_{\mathbf{R}}$, i.e., the $(i_{1},i_{2},i_{3})$-th entry of $\sqrt{\mathcal{W}_{\mathbf{R}}}$ is equal to $\sqrt{(\mathcal{W}_{\mathbf{R}})_{i_{1}i_{2}i_{3}}}$;
  • $\|\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{S}\|_{2,p}$ denotes the weighted tensor $\ell_{2,p}$ (quasi-)norm with $p\in(0,1)$ of a third order tensor $\hat{\mathcal{S}}=\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{S}$ defined by
    (2) $\|\hat{\mathcal{S}}\|_{2,p}=\left(\sum_{i_{3}=1}^{I_{3}}\sum_{i_{2}=1}^{I_{2}}\|\hat{s}_{:i_{2}i_{3}}\|^{p}_{2}\right)^{\frac{1}{p}}=\left(\sum_{i_{3}=1}^{I_{3}}\sum_{i_{2}=1}^{I_{2}}\left(\sum_{i_{1}=1}^{I_{1}}\hat{s}_{i_{1}i_{2}i_{3}}^{2}\right)^{\frac{p}{2}}\right)^{\frac{1}{p}},$
    and the tensor $\ell_{2,p}$ norm is exactly equal to the matrix $\ell_{2,p}$ norm of the unfolding matrix of $\hat{\mathcal{S}}$ along the first dimension, i.e., $\|\hat{\mathcal{S}}\|_{2,p}=\|\hat{S}_{(1)}\|_{2,p}$;
  • $[\mathcal{G}]\times_{1}[X_{1}]\times_{2}[X_{2}]\times_{3}[X_{3}]$ denotes an independent 3-D HOSVD with $[\mathcal{G}]$ being the independent core tensors and $[X_{i}]$ being the independent mode-$i$ factor matrices, $i=1,2,3$, and $[X_{i}]$ is independently orthogonal, i.e., $[{X}_{i}]^{\top}[{X}_{i}]=[\mathbf{I}_{n_{i}}]$, with $\mathbf{I}_{n_{i}}$ representing the identity matrix of size $n_{i}\times n_{i}$;
@@
    with a weight vector $w\in\mathbb{R}^{N}_{++}$.
Model (1) is a nonconvex nonsmooth minimization problem with orthogonal constraints. In particular, the first term of model (1) is a data fidelity term to remove Gaussian noise, the second term is a group sparsity measure to remove sparse noises with linear structures, and the last two terms are the sparsity-enhanced nonlocal low-rank tensor regularization terms. A detailed description of the model for HSI denoising and destriping will be presented in section 5, with the formulations for $\mathbf{R}$, $\mathbf{R}^{\top}$ and $\mathcal{W}_{\mathbf{R}}$.
@@
  • We propose a novel model to remove mixed noise in HSIs using a new sparsity-enhanced low-rank regularization and a tensor $\ell_{2,p}$ norm with $p\in(0,1)$. For removing Gaussian noise, a fourth order tensor is formed by nonlocal similar FBBs via the extraction operator $\mathbf{R}$ and is further regularized on its low-rankness using independent 3-D HOSVD with sparsity enhancement on its core tensors to prompt more low-rankness. The tensor $\ell_{2,p}$ norm for removing the dead lines and stripes is extended from the matrix $\ell_{2,p}$ norm. Some new mathematical results on finding the proximal operator of the tensor $\ell_{2,p}$ norm are also presented. We show that model (1) has a nonempty and bounded solution set.
@@
  • We show the proposed nonlocal low-rank tensor regularized $\ell_{2,p}$ (NLTL2p) approach for HSI denoising and destriping can outperform other state-of-the-art methods including a deep learning based method on the numerical experiments tested on both simulated and real HSI datasets.
The rest of this paper is organized as follows. In section 2, we present some notations and preliminaries for tensors, nonconvex nonsmooth optimization, manifold optimization, and independent 3-D HOSVD. For solving model (1), we propose a P-BCD method in section 3 and conduct its convergence analysis in section 4. Next, we apply the proposed NLTL2p approach to HSI denoising and destriping in section 5 and conduct experiments on simulated and real HSI data in section 6. The concluding remarks are given in section 7.
@@ 2.1 Notations
First, we introduce the notations for tensors. For a third order tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, we let $x_{i_{1}i_{2}i_{3}}$ denote its $(i_{1},i_{2},i_{3})$-th entry, let $x_{:i_{2}i_{3}}$ denote its $(i_{2},i_{3})$-th mode-$1$ fiber and let $X_{::i_{3}}$ denote its $i_{3}$-th frontal slice. And $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}_{++}$ means all its entries are positive. The mode-$k$ unfolding of a third order tensor $\mathcal{X}$ is denoted as $X_{(k)}={\rm unfold}_{(k)}(\mathcal{X})$, which is the process to linearize all indexes except index $k$. The dimensions of $X_{(k)}$ are $I_{k}\times\prod^{3}_{j=1,j\neq k}I_{j}$. An element $x_{i_{1}i_{2}i_{3}}$ of $\mathcal{X}$ corresponds to the position of $(i_{k},j)$ in matrix $X_{(k)}$, where $j=1+\sum^{3}_{l=1,l\neq k}(i_{l}-1)\prod^{l-1}_{m=1,m\neq k}I_{m}$. And the inverse process of the mode-$k$ unfolding of a tensor $\mathcal{X}$ is denoted by $\mathcal{X}={\rm fold}_{(k)}(X_{(k)})$.
@@ 3 Proximal Block Coordinate Descent Algorithm
The proposed nonlocal low-rank tensor model (1) is a nonconvex and nonsmooth optimization problem over Stiefel manifolds. In this section, we propose a P-BCD algorithm for solving model (1).
@@
$\Phi(\mathcal{S},[{X}_{1}],[{X}_{2}],[{X}_{3}],[\mathcal{G}],\mathcal{L}):=$ $\frac{\delta}{2}\|\mathbf{R}(\mathcal{L}+\mathcal{S}-\mathcal{D})\|_{F}^{2}+\gamma\varphi\left(\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{S}\right)+\psi([\mathcal{G}])$
$+\frac{1}{2}H([X_{1}],[X_{2}],[X_{3}],[\mathcal{G}],\mathcal{L}),$
where $\varphi(\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{S})=\|\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{S}\|_{2,p}^{p}$, $\psi([\mathcal{G}])=\|[\mathcal{G}]\|_{1,w}$, and
@@
$\mathcal{S}^{k+1}$ $\in\underset{\mathcal{S}}{\arg\!\min}\;\Phi(\mathcal{S},[{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{\mathcal{S}}}{2}\|\sqrt{\mathcal{W}_{\mathbf{R}}}\odot(\mathcal{S}-\mathcal{S}^{k})\|^{2}_{F},$
@@ 3.1 The update of $\mathcal{S}$
Recall that
$\|\mathbf{R}(\mathcal{L})\|_{F}^{2}=\langle\mathcal{L},\mathbf{R}^{\top}\mathbf{R}(\mathcal{L})\rangle_{F}=\langle\mathcal{L},\mathcal{W}_{\mathbf{R}}\odot\mathcal{L}\rangle_{F}=\|\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{L}\|_{F}^{2}.$
Then $\mathcal{S}^{k+1}$ is computed by
$\mathcal{S}^{k+1}\in\underset{\mathcal{S}}{\arg\!\min}\;\tilde{\gamma}\varphi\left(\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\mathcal{S}\right)+\frac{1}{2}\left\|\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\left[\mathcal{S}-(\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{L}^{k}-\mathcal{D})\right)\right]\right\|_{F}^{2},$
where $\tilde{\gamma}=\frac{\gamma}{\delta+\alpha_{\mathcal{S}}}$ and $\tilde{\alpha}_{\mathcal{S}}=\frac{\delta}{\delta+\alpha_{\mathcal{S}}}$. Rescaling $\mathcal{S}$ using $\sqrt{\mathcal{W}_{\mathbf{R}}}\odot$, $\mathcal{S}^{k+1}$ can be written in terms of the proximal operator of $\varphi$ as follows
(7) $\mathcal{S}^{k+1}\in\left(\sqrt{\mathcal{W}_{\mathbf{R}}}\right)^{-1}\odot\operatorname{prox}_{\tilde{\gamma}\varphi}\left[\sqrt{\mathcal{W}_{\mathbf{R}}}\odot\left(\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{L}^{k}-\mathcal{D}\right)\right)\right],$
where the $(i_{1},i_{2},i_{3})$-th entry of $\left(\sqrt{\mathcal{W}_{\mathbf{R}}}\right)^{-1}$ is equal to $1/\sqrt{(\mathcal{W}_{\mathbf{R}})_{i_{1}i_{2}i_{3}}}$.

### v2

@@ Abstract
Hyperspectral images (HSIs) are often contaminated by a mixture of noises such as Gaussian noise, dead lines, stripes, and so on. In this paper, we propose a novel approach for HSI denoising and destriping, called MLTL2p, which consists of an orthogonal constrained minimization model and an iterative algorithm with convergence guarantees. The model of the proposed MLTL2p approach is built based on a new sparsity-enhanced Multi-scale Low-rank Tensor regularization and a tensor $\ell_{2,p}$ norm with $p\in(0,1)$. The multi-scale low-rank regularization for HSI denoising utilizes the global and local spectral correlation as well as the spatial nonlocal self-similarity priors of HSIs. The corresponding low-rank constraints are formulated based on independent higher-order singular value decomposition with sparsity enhancement on its core tensor to prompt more low-rankness. The tensor $\ell_{2,p}$ norm for HSI destriping is extended from the matrix $\ell_{2,p}$ norm. A proximal block coordinate descent algorithm is proposed in the MLTL2p approach to solve the resulting nonconvex nonsmooth minimization with orthogonal constraints. We show any accumulation point of the sequence generated by the proposed algorithm converges to a first-order stationary point, which is defined using three equalities of substationarity, symmetry, and feasibility for orthogonal constraints. In the numerical experiments, we compare the proposed method with state-of-the-art methods including a deep learning based method, and test the methods on both simulated and real HSI datasets. Our proposed MLTL2p method demonstrates outperformance in terms of metrics such as mean peak signal-to-noise ratio as well as visual quality.
@@ 1 Introduction
Hyperspectral images (HSIs) are collected by hyperspectral sensors across the electromagnetic spectrum. For a three-dimensional (3-D) HSI, the first two dimensions represent spatial information, and the third dimension represents the spectral information of a scene. An illustration of an HSI is shown in Fig. 1. HSIs are widely used for various applications [31, 33, 32, 50] such as object detection [43], material identification [7, 8, 19], etc.
@@
For removing Gaussian noise, the exiting low-rank based methods mainly utilize global spectral correlation of HSIs, and do not use its nonlocal self-similarity prior [5], except for SNLRSF. This limits the performance of restoring clean HSIs. Also, many of those methods use ALM to solve their low-rank regularization based models without providing any convergence analysis. In their ALM, the augmented Lagrangian subproblem is solved by alternating minimization via one Gauss–Seidel-type iteration, instead of an exact joint minimization. As a result, the convergence is not guaranteed for these nonconvex nonsmooth optimization problems with three or more blocks [10].
Recently, some deep neural networks [4, 45, 38] have been proposed to denoise HSIs. Quasi-recurrent neural networks (QRNN) [4] combined recurrent neural networks (RNN) with convolutional neural networks (CNN); the 3-D version of QRNN (QRNN3D) [38] can effectively embed the correlation of HSIs; CNN can also be used as a denoiser in a plug-and-play fashion for HSI denoising [34]. However, the performance of these methods may vary significantly depending on factors such as the datasets used, noise characteristics, and the size of the input data.
To remove Gaussian noise and stripes simultaneously, we propose a new optimization model utilizing low-rank tensor regularization and a group sparsity measure, which is formulated as follows
(1) $\begin{split}\min_{\begin{subarray}{c}\mathcal{S},[{X}_{1}],[{X}_{2}],\\ [{X}_{3}],[\mathcal{G}],\mathcal{L}\end{subarray}}\quad&\frac{1}{2}\|\mathcal{L}+\mathcal{S}-\mathcal{D}\|_{F}^{2}+\gamma\|\mathcal{S}\|^{p}_{2,p}+\|[\mathcal{G}]\|_{1,w}\\ &+\frac{\delta}{2}\|\mathbf{R}(\mathcal{L})-[\mathcal{G}]\times_{1}[X_{1}]\times_{2}[X_{2}]\times_{3}[X_{3}]\|^{2}_{F}\\ \mbox{s.t. }&[{X}_{i}]^{\top}[{X}_{i}]=[\mathbf{I}_{n_{i}}],\quad i=1,2,3,\end{split}$
@@
  • $\|\mathcal{S}\|_{2,p}$ denotes the tensor $\ell_{2,p}$ (quasi-)norm with $p\in(0,1)$ of a third order tensor $\mathcal{S}$ defined by
    (2) $\|\mathcal{S}\|_{2,p}=\left(\sum_{i_{3}=1}^{I_{3}}\sum_{i_{2}=1}^{I_{2}}\|{s}_{:i_{2}i_{3}}\|^{p}_{2}\right)^{\frac{1}{p}}=\left(\sum_{i_{3}=1}^{I_{3}}\sum_{i_{2}=1}^{I_{2}}\left(\sum_{i_{1}=1}^{I_{1}}{s}_{i_{1}i_{2}i_{3}}^{2}\right)^{\frac{p}{2}}\right)^{\frac{1}{p}},$
    and the tensor $\ell_{2,p}$ norm is exactly equal to the matrix $\ell_{2,p}$ norm of the unfolding matrix of ${\mathcal{S}}$ along the first dimension, i.e., $\|{\mathcal{S}}\|_{2,p}=\|{S}_{(1)}\|_{2,p}$;
@@
    with a weight vector $w\in\mathbb{R}^{N}_{++}$;
  • $\mathbf{R}:\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}\to\mathbb{R}^{m_{1}\times m_{2}\times m_{3}\times N}$ denotes the block extraction operator that extracts blocks and forms a fourth order tensor, for example, the nonlocal similar FBB tensor, which can be regularized to exhibit low-rank properties;
  • $[\mathcal{G}]\times_{1}[X_{1}]\times_{2}[X_{2}]\times_{3}[X_{3}]$ denotes an independent 3-D HOSVD with $[\mathcal{G}]$ being the independent core tensors and $[X_{i}]$ being the independent mode-$i$ factor matrices, $i=1,2,3$, and $[X_{i}]$ is independently orthogonal, i.e., $[{X}_{i}]^{\top}[{X}_{i}]=[\mathbf{I}_{n_{i}}]$, with $\mathbf{I}_{n_{i}}$ representing the identity matrix of size $n_{i}\times n_{i}$.
Model (1) is a nonconvex nonsmooth minimization problem with orthogonal constraints. In particular, the first term of model (1) is a data fidelity term to remove Gaussian noise, the second term is a group sparsity measure to remove sparse noises with linear structures, and the last two terms are the sparsity-enhanced nonlocal low-rank tensor regularization terms.
@@
  • We propose a novel model to remove mixed noise in HSIs using a new sparsity-enhanced low-rank regularization and a tensor $\ell_{2,p}$ norm with $p\in(0,1)$. For removing Gaussian noise, a tensor formed via the extraction operator $\mathbf{R}$ is regularized on its low-rankness using independent 3-D HOSVD with sparsity enhancement on its core tensors to prompt more low-rankness. The tensor $\ell_{2,p}$ norm for removing the dead lines and stripes is extended from the matrix $\ell_{2,p}$ norm. Some new mathematical results on finding the proximal operator of the tensor $\ell_{2,p}$ norm are also presented. We show that model (1) has a nonempty and bounded solution set.
@@
  • We propose a multi-scale low-rank tensor regularized $\ell_{2,p}$ (MLTL2p) approach with two phases for HSI denoising and destriping, which utilizes the global and local spectral correlation as well as the spatial nonlocal self-similarity priors of HSIs. We show the MLTL2p approach can outperform other state-of-the-art methods including a deep learning based method on the numerical experiments tested on both simulated and real HSI datasets.
The rest of this paper is organized as follows. In section 2, we present some notations and preliminaries for tensors, nonconvex nonsmooth optimization, manifold optimization, and independent 3-D HOSVD. For solving model (1), we propose a P-BCD method in section 3 and conduct its convergence analysis in section 4. Next, we propose the MLTL2p approach for HSI denoising and destriping in section 5 and conduct experiments on simulated and real HSI data in section 6. The concluding remarks are given in section 7.
@@ 2.1 Notations
First, we introduce the notations for tensors. For a third order tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, we let $x_{i_{1}i_{2}i_{3}}$ denote its $(i_{1},i_{2},i_{3})$-th entry, let $x_{:i_{2}i_{3}}$ denote its $(i_{2},i_{3})$-th mode-$1$ fiber and let $X_{::i_{3}}$ denote its $i_{3}$-th frontal slice. And $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}_{++}$ means all its entries are positive. The mode-$k$ unfolding of a third order tensor $\mathcal{X}$ is denoted as $X_{(k)}={\rm unfold}_{(k)}(\mathcal{X})$, which is the process to linearize all indexes except index $k$. The dimensions of $X_{(k)}$ are $I_{k}\times\prod^{3}_{j=1,j\neq k}I_{j}$. An element $x_{i_{1}i_{2}i_{3}}$ of $\mathcal{X}$ corresponds to the position of $(i_{k},j)$ in matrix $X_{(k)}$, where $j=1+\sum^{3}_{l=1,l\neq k}(i_{l}-1)\prod^{l-1}_{m=1,m\neq k}I_{m}$. The inverse process of the mode-$k$ unfolding of a tensor $\mathcal{X}$ is denoted by $\mathcal{X}={\rm fold}_{(k)}(X_{(k)})$. An illustration of the unfoldings is presented in Fig. 2.
Figure 2: The unfoldings of tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$
@@ 2.2 Independent 3-D HOSVD
Figure 3: A stack of independent third order tensors
Also, a stack of independent matrices is denoted as $[X]\in\mathbb{R}^{m\times n\times N}$. And we say $[X]$ is independently orthogonal if $[{X}]^{\top}[{X}]=[\mathbf{I}_{n}]$, meaning $([{X}]^{(j)})^{\top}[{X}]^{(j)}=\mathbf{I}_{n}$, or equivalently $[X]\in[\mathbb{S}_{m,n}]$, meaning $[X]^{(j)}\in\mathbb{S}_{m,n}$. Then we define an independent (truncated) 3-D HOSVD of $[\mathcal{Y}]$ as
@@ 3 Proximal Block Coordinate Descent Algorithm
The proposed low-rank tensor model (1) is a nonconvex and nonsmooth optimization problem over Stiefel manifolds. In this section, we propose a P-BCD algorithm for solving model (1).
@@
$\Phi(\mathcal{S},[{X}_{1}],[{X}_{2}],[{X}_{3}],[\mathcal{G}],\mathcal{L}):=$ $\frac{1}{2}\|\mathcal{L}+\mathcal{S}-\mathcal{D}\|_{F}^{2}+\gamma\varphi\left(\mathcal{S}\right)+\psi([\mathcal{G}])$
$+\frac{\delta}{2}H([X_{1}],[X_{2}],[X_{3}],[\mathcal{G}],\mathcal{L}),$
where $\varphi(\mathcal{S})=\|\mathcal{S}\|_{2,p}^{p}$, $\psi([\mathcal{G}])=\|[\mathcal{G}]\|_{1,w}$, and
@@
$\mathcal{S}^{k+1}$ $\in\underset{\mathcal{S}}{\arg\!\min}\;\Phi(\mathcal{S},[{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})+\frac{\alpha_{\mathcal{S}}}{2}\|\mathcal{S}-\mathcal{S}^{k}\|^{2}_{F},$
@@ 3.1 The update of $\mathcal{S}$
We update $\mathcal{S}^{k+1}$ using the proximal operator of $\varphi$ as follows
$\mathcal{S}^{k+1}\in$ $\underset{\mathcal{S}}{\arg\!\min}\;\tilde{\gamma}\varphi\left(\mathcal{S}\right)+\frac{1}{2}\left\|\mathcal{S}-(\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{L}^{k}-\mathcal{D})\right)\right\|_{F}^{2},$
(7) $=$ $\operatorname{prox}_{\tilde{\gamma}\varphi}\left(\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{L}^{k}-\mathcal{D}\right)\right),$
where $\tilde{\gamma}=\frac{\gamma}{1+\alpha_{\mathcal{S}}}$ and $\tilde{\alpha}_{\mathcal{S}}=\frac{1}{1+\alpha_{\mathcal{S}}}$.